Leonardo, who is married but files separately, earns $80,000 of taxable income. He also has $15,000 in city of Tulsa bonds. His wife, Theresa, earns $50,000 of taxable income.

If Leonardo and his wife file married filing jointly in 2019, What is Leonardo and Theresa's effective tax rate for 2019 (rounded)? (Use tax rate schedule.)

15.63 percent

22.00 percent

12.00 percent

21.04 percent

None of the choices are correct
Answer #1
Total taxable income 130000 =80000+50000
Tax on $130000 20317 =9086+(130000-78950)*22%
Tax payable 20317
Divide by Total Income 145000 =130000+15000
Effective tax rate for 2019 14.01%
Option E None of the choices are correct is correct option